# Bouncing cosmology and dynamical analysis of stability with non-minimal kinetic coupled gravity

### v2

@@ Abstract
In this paper, we model the bounce phase, stability and the reconstruction of the universe by non-minimal kinetic coupling. In the process, we obtained importance information about the energy density and the matter pressure of the universe in relation to the previous universe through the bounce quantum phase. The novelty of the work is that the scale factor is obtained directly from the model and is fitted with an exponential function, with this view we explore the process of the early universe even the bounce phase. After that, we plot the cosmological parameters in terms of time evolution. In what follows, we investigate the stability of the model by the dynamical system analysis in a phase plane. Finally, we examine the stability of the universe, especially in the inflationary period, by using the phase space trajectories.
@@ I Introduction
The importance of modified gravity theories in the study of the very universe is illustrated by the literature review of Odintsov et al. Odintsov-2023 . The review covers many aspects of modified gravity in cosmology and concentrated on the recent advances on inflationary cosmology in the modified gravity frame-work. It turns out that inflation as supported by theory and observation, is the fundamental aspects of the evolutionary characteristic of the universe, and that modified theories of gravity play a central role in this aspect because they describe able to describe a inflation field as multi fold due to its multiple short comings. Among the modified gravity theories, non-minimal gravitational theory enable one to obtain the asymmetric bounce, and do the dynamical stability analysis Banerjee et al and Odintsov et al. Banerjee-2022 ; Odintsov-2021 ; Odintsov-2020 . More relevantly to our work in this paper, Kerachian et al. Kerachian-2019 analyzed dynamics of a non-minimally coupled scalar field with an unspecified positive potential in a spatially curved FRW spacetime, and studied; its dynamics in a spatially curved FRW spacetime, defined a set of dimensionless variables which are valid for all positive, negative, and zero curvatures. The crucial point being able to define the dimensionless variables for each physical quantity so that a better interpretation of the full dynamics of the system can be attained. Further more the critical points were derived for a generic potential and singularities of the system, and then demonstrated that the determined critical points actually correspond to the de Sitter universe, the radiation universe, and the Milne universe.
@@
Since the scale factor can never be zero, then the space-time singularity does not occur in the early universe, as a result, the bounce scenario becomes stronger, i.e., when the scale factor reaches its minimum size, the universe begins to grow. This expression mathematically means that the slope of the scale factor or the derivative of the scale factor is zero at the mentioned point, so we call this point the bounce point. In fact, before reaching the bounce point, the previous universe contracts, and after passing this point, our new universe begins to grow again. While the size of the scale factor is decreasing at the end of the previous universe period, we can say $\dot{a}<0$, and while the size of the scale factor is increasing at the beginning of the current universe, we can say $\dot{a}>0$. From what was said, it can be concluded that the bounce occurs when the Hubble parameter is zero, and for the collapse of the previous universe and the beginning of the current universe is $H<0$ and $H>0$, respectively Cai-2014 ; Brandenberger-2012 ; Cai-2011 ; Sadeghi-2009 ; Singh-2018 ; Caruana-2020 ; Minas-2019 ; Skugoreva-2020 ; Amani-2016 ; Nojiri-2016 ; Cai-2013 ; Battefeld-2015 ; Sadeghi-2010 ; Lohakare-2022 ; Ilyas-2021 ; Ijjas-2018 ; Myrzakulov-2014 ; Shabani-2018 ; Brandenberg-2017 ; Bamba-2014 ; Tripathy-2019 ; Nojiri-2017 ; Brizuela-2010 ; Martin-Benito-2021 ; Sato-2018 ; Piao-2005 ; Cai-2009 ; Gordon-2003 ; Chattopadhyay-2017 ; Kerlick-1976 ; Alberti-2020 ; Novello-2008 ; Makarenko-2017 ; Brandenberger-2012a ; Cai-2013a ; Agrawal-2023 .
@@ III Bouncing cosmology behavior
where the energy density is $\rho>0$, and therefore the EoS is less than $-1$ ($\omega<-1$).
@@
where by applying the bouncing condition (10) we will have
$\dot{\phi}_{b}^{2}F_{b}>\frac{1}{\kappa^{2}\xi},~{}~{}~{}~{}\dot{H}_{b}F_{b}>\frac{1}{2\xi},$ (12)
@@
$V(\phi)=\frac{\textrm{v}_{0}}{e^{\lambda\phi}+e^{-\lambda\phi}}$ (13)
@@
where $\textrm{v}_{0}$ and $\lambda$ are constant. This type of potential ($\sim\textrm{sech}(\lambda\phi)$), which is associated with scalar fields, determines how the field evolves over time and how it affects the early expansion of the universe. In this case, the corresponding potential shape implies a scalar field that can roll down from the hilltop of near the origin and end up in the inflation period after passing through a steep region. Therefore, it could have implications for the production of perturbations in the early universe and the subsequent formation of large-scale structures. In what follows, considering the effect of the third and fourth terms of action (3) on the corresponding potential, it is necessary to consider the form of function $F(\phi)$ as a power law to make a balance between potential and non-minimal kinetic gravity. However, we consider $F(\phi)=\sum\limits_{n=0}c_{n}\phi^{2n}$ in which $c_{0}=1$, $c_{1}=c$, and $c_{n\geq 2}=0$.
@@
Now, according to what was said above we have to solve the corresponding system. Since the solution of the present model has some complications, then we turn to the numerical solution. For this purpose, we numerically solve the system of Eqs. (7a) and (7c) for two functions of the scale factor and the scalar field in terms of cosmic time. On the other hand, the free parameters play a very important role, hence their selection is very sensitive according to the expected results. These free parameters must be chosen in such a way that the variation of the Hubble parameter in terms of cosmic time shows a transition from the negative value to the positive value with a positive slope, i.e., the bouncing solution. Therefore, we plot the graphs of cosmological parameters in terms of cosmic time with the approved values of free parameters as $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$. Fig. 1 (the solid graph) shows us that we have $\dot{a}<0$ when the universe is in the pre-big bang or contraction phase, and $\dot{a}>0$ when the universe is in the post-big bang or expansion phase. This means that the scale factor has a negative slope before the big bang and a positive slope for the expansion phase. It should be noted that the result obtained is from a numerical solution and not from an analytical solution, so it is interesting to be able to express a purely mathematical function for the scale factor in the following form
@@
Figure 1: Graphs of the scale factor in terms of time evolution in the form of the numerical solution and the adopted scale factor.
Another cosmological parameters that can describe the bouncing solution is the Hubble parameter. For this purpose, we can draw the Hubble parameter from the scale factor in terms of cosmic time, with the help of the mentioned numerical solution as a solid graph shown in Fig. 2. Therefore, Fig. 2 shows us that the Hubble parameter passes from the contraction phase, $H<0$, to the expansion phase, $H>0$, with a positive slope, and we have the bouncing point ($H=0$) between two phases. In other words, from the previous universe, the value of the Hubble parameter decreases exponentially, and then when the universe approaches its bounce point it start to negatively increase exponentially till it reach zero at the bounce point (still at quantum gravity), and there after increase exponentially during the inflationary phase, and then start decrease exponentially in power tail after the inflation. Therefore, we can obtain the calculated Hubble parameter from the adopted scale factor (14) as an explicit function in terms of cosmic time in the following form
@@
where the corresponding coefficients are the same as the obtained values by fitting them to the adopted scale factor.
@@
Figure 2: Graphs of the Hubble parameter in terms of time evolution in the form of the numerical solution and the calculated scale factor.
@@
In what follows, we plot the scalar field in terms of time evolution with the same values of free parameters $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$, as shown in Fig. 3. Since scalar fields play an important role in cosmology, it is therefore very efficient for calculating cosmological parameters. We note that the effect of the scalar field is directly used as an auxiliary field in the calculation of cosmological parameters. Therefore, only the graph of the scalar field in terms of time evolution shown in Fig. 3 is sufficient, so that, there is no need to obtain the analytical form of its function.
@@
Figure 3: Graph of the scalar field in terms of time evolution by $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
@@
Next, we explore other cosmological parameters, including energy density and pressure around the bouncing point. In order to continue the numerical solution of the current system, we draw the variation of energy density and pressure resulting from Eqs. (8) in terms of time evolution as shown in Figs. 4. As it is evident from the energy density and the pressure of matter in Fig. 4, we note that the present universe entered the unstable quantum gravity phase at the bounce phase, from the stable later stage of the previous universe. Then it proceeded through an exponentially expanded phase (inflationary phase) till at about $t=5$, when it flatten out once more as in the previous universe. The pressure of the matter content of the previous universe at a late stage seems to be negatively increasing till at the bounce between $t=-5$ and $t=5$ when pressure gradient changes from negative to positive. From $t=5$ the universe seems to be decreasing quite rapidly and then it started decreasing asymptotically at about $t=10$.
@@
Figure 4: Graphs of the energy density and the pressure in terms of time evolution by $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
@@ V Conclusion
In this paper, we have modeled the evolution of the universe from the point of view of bouncing cosmology using the theory of non-minimal kinetic. For this purpose, we first addressed the issue of non-minimum kinetic theory in background flat-FRW. We restricted the minimally coupled terms $\xi$ and $\eta$ to the condition $\eta+2\xi=0$ to modify the corresponding action from the more general case to a more well-known case of the Einstein tensor. Next, we obtained the modified forms of Friedmann’s equations and the field equation, and then we acquired the energy density and the pressure of the universe.
@@
In what follows, to investigate bouncing cosmology, we obtained the bouncing condition and we then reconstructed the evolution of the universe using the theory of non-minimal kinetic coupling. By applying the bouncing condition in the Friedman equations and the field equation, we calculated the scale factor and the Hubble parameter in terms of time evolution with numerical solution. We earned the purely mathematical function for the scale factor in the form of a exponentially function, so that the adopted scale factor was fitted with the scale factor graph points resulting from the numerical solution. In that case, we also obtained a mathematical function for the Hubble parameter from the adopted scale factor. We can see from the plot of the Hubble parameter that it passes through a contraction phase ($H<0$), to a bounce point ($H=0$) and then into an expansion phase ($H>0$), with a positive slope. Afterward, the graphs of the scalar field, the energy density, and the pressure in terms of cosmic time were plotted.
@@
To explore the stability of the model, we used dynamical system analysis in phase plane by critical or fixed points. For the same purpose, we introduced the dimensionless variables as an autonomous dynamical system, and then we earned the autonomous equations of the cosmic dynamic system. Afterward, by setting the autonomous equations equal to zero, we obtained the fixed points according to Tab. 1, next, stability conditions of critical points were obtained according to the contents of Tab. IV. In what follows, we considered four different modes for the fixed points and then, obtained the stability points by placing constraints for each mode. After that, we drew the figures of the phase space trajectories in $x-y$ phase plane. The results obtained from these modes showed that, after the bounce point, the universe enters different periods, one of which is cosmic inflation. Since the shape of the potential has a direct effect on the stability in the inflationary period, so, choosing potential (13) brought good results for the inflationary period. Finally, the obtained results help us to better understand in the early universe. As an interesting future work, we propose the extension of the present research for a viable model.

### v3

@@ Abstract
In this paper, we model the bounce phase, stability, and the reconstruction of the universe by non-minimal kinetic coupling. In the process, we obtained importance information about the energy density and the matter pressure of the universe in relation to the previous universe through the bounce quantum phase. The novelty of the work is that the scale factor is obtained directly from the model and is fitted with an exponential function, with this view we explore the process of the early universe even the bounce phase. After that, we plot the cosmological parameters in terms of time evolution. In what follows, we investigate the stability of the model by dynamical system analysis in a phase plane. Finally, we examine the stability of the universe, especially in the inflationary period, by using the phase-space trajectories.
@@ I Introduction
The importance of modified gravity theories in the study of the very universe is illustrated by the literature review of Odintsov et al. Odintsov-2023 . The review covers many aspects of modified gravity in cosmology and is focused on recent advances on inflationary cosmology in the modified gravity framework. It turns out that inflation as supported by theory and observation is the fundamental aspects of the evolutionary characteristic of the universe, and that modified theories of gravity play a central role in this aspect because they describe able to describe an inflation field as multi-fold due to its multiple shortcomings. Among the modified gravity theories, non-minimal gravitational theory enables one to obtain the asymmetric bounce, and perform the dynamical stability analysis Banerjee and et al, and Odintsov and et al. Banerjee-2022 ; Odintsov-2021 ; Odintsov-2020 . More relevantly to our work in this paper, Kerachian et al. Kerachian-2019 analyzed the dynamics of a non-minimally coupled scalar field with an unspecified positive potential in a spatially curved FRW spacetime, and studied; its dynamics in a spatially curved FRW spacetime, defined a set of dimensionless variables which are valid for all positive, negative, and zero curvatures. The crucial point being able to define the dimensionless variables for each physical quantity so that a better interpretation of the full dynamics of the system can be attained. Furthermore the critical points were derived for a generic potential and singularities of the system, and then demonstrated that the determined critical points actually correspond to the de Sitter universe, the radiation universe, and the Milne universe.
@@
Since the scale factor can never be zero, then the space-time singularity does not occur in the early universe, as a result, the bounce scenario becomes stronger, i.e., when the scale factor reaches its minimum size, the universe begins to grow. This expression mathematically means that the slope of the scale factor or the derivative of the scale factor is zero at the mentioned point, so we call this point the bounce point. In fact, before reaching the bounce point, the previous universe contracts, and after passing this point, our new universe begins to grow again. While the size of the scale factor is decreasing at the end of the previous universe period, we can say $\dot{a}<0$, and while the size of the scale factor is increasing at the beginning of the current universe, we can say $\dot{a}>0$. From what was said, it can be concluded that the bounce occurs when the Hubble parameter is zero, and for the collapse of the previous universe and the beginning of the current universe is $H<0$ and $H>0$, respectively Cai-2014 ; Brandenberger-2012 ; Cai-2011 ; Sadeghi-2009 ; Singh-2018 ; Caruana-2020 ; Minas-2019 ; Skugoreva-2020 ; Amani-2016 ; Nojiri-2016 ; Cai-2013 ; Battefeld-2015 ; Sadeghi-2010 ; Lohakare-2022 ; Ilyas-2021 ; Ijjas-2018 ; Myrzakulov-2014 ; Shabani-2018 ; Brandenberg-2017 ; Bamba-2014 ; Tripathy-2019 ; Nojiri-2017 ; Brizuela-2010 ; Martin-Benito-2021 ; Sato-2018 ; Piao-2005 ; Cai-2009 ; Gordon-2003 ; Chattopadhyay-2017 ; Kerlick-1976 ; Alberti-2020 ; Novello-2008 ; Makarenko-2017 ; Brandenberger-2012a ; Cai-2013a ; Agrawal-2023 ; Sadatian-2024 ; Battista-2021 ; Assolohou-2012 .
@@ III Bouncing cosmology behavior
where $\rho+p<0$, and this is the same statement regarding the violation of the null energy condition. So this violation helps the universe to continue contracting before the Big Bang and not go back, so that the cosmic bounce occurs and the universe continues to expand after the Big Bang.
@@
where by applying the bouncing condition (10), and violation of the null energy condition ($\rho+p<0$) we will have
$\xi\dot{H}_{b}F_{b}>\frac{1}{2},$ (12)
@@
$V(\phi)=\frac{\textrm{v}_{0}}{\cosh({\lambda\phi})},$ (13)
@@
where $\textrm{v}_{0}$ and $\lambda$ are constant. This potential has been studied in many cosmological models, including scalar fields, tachyon fields, teleparallel gravity, and brane–world gravity Sahni-2000 ; Singh-2003 ; Geng-2012 ; Steer-2004 ; Leblond-2003 . This type of potential, which is associated with scalar fields, determines how the field evolves over time and how it affects the expansion of the early universe. In this case, the corresponding potential shape implies a scalar field that can roll down from the hilltop of near the origin and end up in the inflation period after passing through a steep region. Therefore, it could have implications for the production of perturbations in the early universe and the subsequent formation of large-scale structures. In what follows, considering the effect of the third and fourth terms of action (3) on the corresponding potential, it is necessary to consider the form of function $F(\phi)$ as a power law to make a balance between potential and non-minimal kinetic gravity. However, we consider $F(\phi)=\sum\limits_{n=0}c_{n}\phi^{2n}$ in which $c_{0}=1$, $c_{1}=c$, and $c_{n\geq 2}=0$.
@@
Now, according to what was said above we have to solve the corresponding system. Since the solution of the present model has some complications, then we turn to the numerical solution. For this purpose, we numerically solve the system of Eqs. (7a) and (7c) for two functions of the scale factor and the scalar field in terms of cosmic time. On the other hand, the free parameters play a very important role, hence their selection is very sensitive according to the expected results. These free parameters must be chosen in such a way that the variation of the Hubble parameter in terms of cosmic time shows a transition from the negative value to the positive value with a positive slope, i.e., the bouncing solution. So, in order to plot the cosmological parameters, it is necessary to know the free parameter values. Therefore, we can consider a set of free parameter values taking into account the above constraints. With this in mind, we are limited to only a series of free parameter values, namely $\xi=20$, $\textrm{v}_{0}=12.5$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$, to examine the present model. Fig. 1 shows us that we have $\dot{a}<0$ when the universe is in the pre-big bang or contraction phase, and $\dot{a}>0$ when the universe is in the post-big bang or expansion phase. This means that the scale factor has a negative slope before the big bang and a positive slope for the expansion phase. The advantage of this research is that due to the non-minimum kinetic term, we see an asymmetric cyclic universe, so that the collapse of the pre-Big Bang inflationary universe is faster than the evolution of the post-Big Bang inflationary universe. This is because it involves fundamental changes in the state of space-time and matter at that epochs. Therefore, the change from before the Big Bang to after it is not physically and theoretically symmetric, but rather a kind of unique bouncing and evolution in the state of the universe.
Figure 1: Graph of the scale factor in terms of cosmic time.
Another cosmological parameters that can describe the bouncing solution is the Hubble parameter. For this purpose, we can draw the Hubble parameter in terms of cosmic time as shown in Fig. 2 with the numerical solution of the scale factor (Fig. 1). Therefore, Fig. 2 shows us that the Hubble parameter passes from the contraction phase, $H<0$, to the expansion phase, $H>0$, with a positive slope, and we have the bouncing point ($H=0$) between two phases. In other words, from the previous universe, the value of the Hubble parameter decreases exponentially, and then when the universe approaches its bounce point it begins to negatively increase exponentially till it reach zero at the bounce point (still at quantum gravity), and there after increase exponentially during the inflationary phase, and then start decrease exponentially in power tail after the inflation.
Figure 2: Graph of the Hubble parameter in terms of cosmic time.
One of the important tools for identifying and analyzing the early turbulent period in bouncing cosmology is to study the behavior of quantity $\frac{1}{aH}$. The behavior of this quantity during different stages (contraction, bounce, and expansion) provides information about how turbulence grows and evolves. This analysis improves our understanding of the early stages of the universe, so we plot the variation of $\frac{1}{aH}$ with respect to cosmic time as shown in Fig. 3. The important features in graph 3 are as follows:
  • Before the Bounce ($t<0$): $\frac{1}{aH}$ decreases as the universe contracts.
  • At the Bounce ($t=0$): The Hubble parameter $H$ approaches zero, causing an instantaneous divergence of $\frac{1}{aH}$.
  • After the Bounce ($t>0$): $\frac{1}{aH}$ increases as the universe expands.
From this pattern, we can see that our universe is transitioning from a state of contraction to expansion.
Figure 3: Graph of $\frac{1}{aH}$ in terms of cosmic time.
Now, we want to fit the numerical solution of Fig. 1 by an exponential function that is an appropriate descriptor based on the principles of the early universe. Therefore, the following exponential function is introduced for the scale factor in the form
@@
Figure 4: Graph of the scale factor in terms of cosmic time in the form of the numerical solution and the adopted scale factor.
Therefore, we can obtain the calculated Hubble parameter from ASF (14) as an explicit function in terms of cosmic time in the following form
@@
where the corresponding coefficients are the same as the obtained values by fitting them to ASF. In that case, by substituting the above fitted values $A$, $B$, $C$, $D$, and $\mu$ into Eq. (15), we can plot both the ”calculated Hubble parameter” and ”obtained numerical solution” graphs as shown in Fig. 5. Note that the ”calculated Hubble parameter” obtains from ASF of Eq. (14).
@@
Figure 5: Graph of the Hubble parameter in terms of cosmic time in the form of the numerical solution and the calculated Hubble parameter.
@@
In what follows, we plot the scalar field in terms of cosmic time numerically with the same values of free parameters $\xi=20$, $\textrm{v}_{0}=12.5$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$, as shown in Fig. 6. Since scalar field plays an important role in cosmology, it is very efficient for calculating cosmological parameters. We note that the effect of the scalar field is directly used as an auxiliary field in the calculation of cosmological parameters. Therefore, only the graph of the scalar field in terms of cosmic time shown in Fig. 6 is sufficient, so that, there is no need to obtain the analytical form of its function.
@@
Figure 6: Graph of the scalar field in terms of cosmic time by $\xi=20$, $\textrm{v}_{0}=12.5$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
@@
Next, we explore other cosmological parameters, including energy density and pressure around the bouncing point. In order to continue the numerical solution of the current system, we draw the variation of energy density and pressure resulting from Eqs. (8) in terms of cosmic time as shown in Figs. 7. As it is evident from the energy density and the pressure of matter in Fig. 7, the present universe entered the unstable quantum gravity phase at the bounce phase, from the stable later stage of the previous universe. From the variation of energy density in Fig. 7, we see that the universe entered an inflationary period immediately after the Big Bang. In that case, the energy density suddenly increased, and after the end of this period, with the expansion of the universe, the variation of in energy density gradually decreased because the expansion of the universe causes the expansion of space and the conversion of energy into matter and vice versa. Therefore, fluctuations in energy density are transmitted from the pre-Big Bang to the post-Big Bang phase through mechanisms such as gravitational amplification, spectral transfer of fluctuations, suppression of anomalies, and tuning by scalar fields. These fluctuations appear in the form of large-scale structures, gravitational waves, and cosmic microwave background radiation, providing us with valuable information about phase transitions and early processes in the universe. On the other hand, the variation of pressure proceeded through an exponentially expanded phase (inflationary phase) till at about $t=5$, when it flatten out once more as in the previous universe. The pressure of the matter content of the previous universe at a late stage seems to be negatively increasing till at the bounce between $t=-5$ and $t=5$ when pressure gradient changes from negative to positive. From $t=5$ the universe seems to be decreasing quite rapidly and then it started decreasing asymptotically at about $t=10$.
@@
Figure 7: Graphs of the energy density and the pressure in terms of time evolution by $\xi=20$, $\textrm{v}_{0}=12.5$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
Another fundamental parameter of cosmology is the investigation of the EoS parameter. Therefore, following the numerical solution above, we plot the EoS parameter in terms of cosmic time, as shown in Fig. 8. The trajectory representation in Fig. 8 shows us that the universe moves from the phantom region to the quintessence region in the pre-Big Bang, and then enters the post-Big Bang stage after passing the bouncing point. At this time, the universe begins to grow from the quintessence region and, after passing through the early period of the universe, enters the phantom region. Therefore, the results obtained can be a good candidate as viable cosmological model from the early universe to the late universe.
Figure 8: Graph of the EoS parameter in terms of cosmic time by $\xi=20$, $\textrm{v}_{0}=12.5$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
@@ V Conclusion
In this paper, we have modeled the evolution of the universe from the point of view of bouncing cosmology within the framework of non-minimal kinetic theory. For this purpose, we first addressed the issue of non-minimum kinetic theory in background flat-FRW. In this way, we restricted the minimally coupled terms $\xi$ and $\eta$ to the condition $\eta+2\xi=0$ to modify the corresponding action from the more general case to a more well-known case of the Einstein tensor. Next, we obtained the modified forms of Friedmann’s equations and the field equation, and then we acquired the energy density and the pressure of the universe.
@@
In what follows, to investigate bouncing cosmology, we obtained the bouncing condition and we then reconstructed the evolution of the universe using non-minimal kinetic coupling theory. By applying the bouncing condition in the Friedman equations and the field equation, we calculated the scale factor and the Hubble parameter in terms of cosmic time with numerical solution. We earned the purely mathematical function for the scale factor in the form of a exponentially function (14), so that ASF was fitted with the scale factor graph points resulting from the numerical solution. In that case, we also obtained a mathematical function for the Hubble parameter (15) from ASF. We can see from the graph of the Hubble parameter that it passes through a contraction phase ($H<0$), to a bounce point ($H=0$) and then into an expansion phase ($H>0$), with a positive slope. Afterward, the graphs of the scalar field, the energy density, the pressure, and the EoS parameter in terms of cosmic time were plotted. These graphs have shown the variations from the pre-Big Bang to the post-Big Bang state. It has exclusively shown the variation of the EoS parameter in terms of cosmic time from the early universe to the late universe, i.e., evolution of the universe has been explained as a viable cosmological model.
@@
To explore the stability of the model, we used dynamical system analysis in phase plane by critical or fixed points. For the same purpose, we introduced the dimensionless variables as an autonomous dynamical system, and then we earned the autonomous equations of the cosmic dynamic system. Afterward, by setting the autonomous equations equal to zero, we obtained the fixed points according to Tab. 1, next, stability conditions of critical points were obtained according to the contents of Tab. IV. In what follows, we considered four different modes for the fixed points and then, obtained the stability points by placing constraints for each mode. After that, we drew the figures of the phase space trajectories in $x-y$ phase plane. The results obtained from these modes showed that, after the bounce point, the universe enters different periods, one of which is cosmic inflation. Since the shape of the potential has a direct effect on the stability in the inflationary period, so, choosing potential (13) brought good results for the inflationary period. Finally, the results obtained help us to better understand the early universe and even the evolution of the universe as a viable model.